# A Theoretical Framework for Grokking: Interpolation followed by Riemannian Norm Minimisation

### v1

@@ 4 Grokking as two-timescale dynamics
This section states our main results, where we characterise the two-timescale dynamics of the regularised gradient flow (1) in the limit $\lambda\to 0$. In Section 4.1 we describe the first phase, the fast dynamics, where $w^{\lambda}$ approximates the unregularised gradient flow solution on finite time horizons. Section 4.2 then identifies the second, slow dynamics happening at arbitrarily large time horizons, where $w^{\lambda}$ follows the Riemannian flow of the $\ell_{2}$ norm on the manifold $\mathcal{M}$ of stationary points.
@@ 4.1 Fast dynamics
A first simple observation is that as $\lambda\to 0$, $F_{\lambda}\to F$ uniformly on any compact of $\mathbb{R}^{d}$. From there, it seems natural that $w^{\lambda}$ should converge, at least pointwise, to $w^{\mathrm{GF}}$ as $\lambda\to 0$. A Grönwall argument indeed allows to characterise the first, fast timescale dynamics given by Proposition 1 below.
@@
If Assumption 1 holds, then for all $T\geq 0$, $w^{\lambda}\underset{\lambda\to 0}{\longrightarrow}w^{GF}$ in $(\mathcal{C}^{0}([0,T],\mathbb{R}^{d}),\|\cdot\|_{\infty})$
Importantly, uniform convergence only holds on finite time intervals of the form $[0,T]$, but is not true on $\mathbb{R}_{+}$. More precisely, grokking is observed when the two limits cannot be exchanged: $\lim_{\lambda\to 0^{+}}\lim_{t\to\infty}w^{\lambda}(t)\neq\lim_{t\to\infty}\lim_{\lambda\to 0^{+}}w^{\lambda}(t)=\lim_{t\to\infty}w^{\mathrm{GF}}(t)$. In Figure 1, the endpoint of the red arrow corresponds to this first limit, while the endpoint of the blue arrow corresponds to the second. This distinction highlights a key aspect of grokking: the dynamics evolve on two different timescales. Initially, as described by Proposition 1, the regularised flow tracks the unregularised one. But at much larger time horizons, the regularised dynamics begin to diverge.
@@ 4.2 Slow dynamics
$\phi(w,t)=\phi(w,0)-\int_{0}^{t}\nabla F(\phi(w,s))\ \mathrm{d}s,\qquad\forall(w,t)\in\mathbb{R}^{d}\times\mathbb{R}_{+}.$
Note that $\phi(w,t)$ simply corresponds to the solution of the gradient flow of Equation 2 at time $t$ when initialised in $w$. When possible to define—i.e., when the gradient flow admits a limit point in $\mathbb{R}^{d}$—we define the mapping $\Phi$ as
@@
Thanks to Assumption 1, the unregularised flow initialised in $w_{0}$ admits the limit point $\Phi(w_{0})$, which is necessarily a stationary point of the loss, i.e., $\nabla F(\Phi(w_{0}))=0$. Assumption 2 then ensure that the mapping $\Phi$ is defined and $\mathcal{C}^{2}$ on some neighbourhood of $\mathcal{M}$, thanks to a result of Falconer (1983).
@@
If Assumptions 1 and 2 hold, there exists an open neighbourhood $U$ of $\mathcal{M}$ such that $\Phi$ is defined and $\mathcal{C}^{2}$ on $U$.
@@
Our goal in this section is to study the limit function $\lim_{\lambda\to 0}\tilde{w}^{\lambda}$. Intuitively, the $\frac{1}{\lambda}\nabla F$ term in Equation 5 will enforce this limit to stay on the stationary manifold $\mathcal{M}$ for any $t>0$. The slow dynamics will be shown to approximate the Riemannian flow of the squared Euclidean norm on the stationary manifold $\mathcal{M}$ for $t>0$. This limit flow is defined by $\tilde{w}^{\circ}$, which is the solution of the following differential equation on $\mathbb{R}_{+}$, for the function $\ell_{2}:w\mapsto\|w\|_{2}^{2}/2$,
@@
Recall that the Riemannian gradient is $\mathrm{grad}_{\mathcal{M}}\ \ell_{2}(w)=P_{\mathrm{Ker}(\nabla^{2}F(w))}(w)$ for any $w\in\mathcal{M}$. Denoting $D\Phi_{w}$ the differential of $\Phi$ at $w$, Li et al. (2021, Lemma 4.3) proved that for any $w\in\mathcal{M}$, $P_{\mathrm{Ker}(\nabla^{2}F(w))}=D\Phi_{w}$, i.e., the differential of $\Phi$ at $w$ is given by the orthogonal projection onto the kernel space of the Hessian of $F$. In consequence, $\tilde{w}^{\circ}$ also satisfies the following differential equation:
@@
Using this alternative description of $\tilde{w}^{\circ}$, we can now prove our main result, given by Proposition 2.
@@
If Assumptions 1 and 2 hold, then for all $T,\varepsilon>0$, we have $\tilde{w}^{\lambda}\underset{\lambda\to 0}{\longrightarrow}\tilde{w}^{\circ}$ in $(\mathcal{C}^{0}([\varepsilon,T],\mathbb{R}^{d}),\|\cdot\|_{\infty})$, where $\tilde{w}^{\circ}$ is the unique solution on $\mathbb{R}_{+}$ of the differential equation (6).
Proposition 2 states that the slow dynamics $\tilde{w}^{\lambda}$ converges uniformly to $\tilde{w}^{\circ}$ as $\lambda\to 0$ on any compact interval of the form $[\varepsilon,T]$. Note that excluding 00 from this interval (i.e., $\varepsilon>0$) is necessary. Indeed, uniform convergence cannot happen on an interval of the form $(0,T]$, since $\tilde{w}^{\lambda}(0)=w_{0}$ for any $\lambda>0$ and $\tilde{w}^{\circ}(0)=\Phi(w_{0})$. In particular, Proposition 2 leads to pointwise convergence of $\tilde{w}^{\lambda}$: we have
@@
This limit function $\lim_{\lambda\to 0}\tilde{w}^{\lambda}$ is non-continuous at 00. Indeed the whole fast dynamics, which follows the unregularised flow, happens at that 00 point in the limit $\lambda\to 0$. On the other hand, Proposition 2 describes the second phase of the dynamics, starting from the convergence point of the unregularized flow $\Phi(w_{0})$ – at the rescaled time $0^{+}$ – and following the Riemannian flow on $\mathcal{M}$.
Note that, once the junction between the slow and fast dynamics is carefully handled via Lemma 2 in Section A.2, Proposition 2 can be derived from Fatkullin et al. (2010, Theorem 2.2), which heavily relies on the technical result of Katzenberger (1990). However, for the sake of completeness and readability, we provide a concise and self-contained proof of Proposition 2, avoiding the use of heavyweight methods from Katzenberger (1990) and relying on weaker assumptions.
@@
We here provide a sketch of proof with the key arguments leading to Proposition 2. Its complete and detailed proof can be found in Section A.3. We first define the shifted slow dynamics $\tilde{v}^{\lambda}$ for any $t\geq 0$ as $\tilde{v}^{\lambda}(t)=\tilde{w}^{\lambda}(t+t(\lambda))$, where $t(\lambda)$ is the “junction point” between the two dynamics given by Lemma 2 in Section A.2, and satisfies $\lim_{\lambda\rightarrow 0}t(\lambda)=0$. Using Lemma 2, $\tilde{v}^{\lambda}$ then follows the same differential equation as $\tilde{w}^{\lambda}$, with an initial condition now satisfying $\lim_{\lambda\to 0}\tilde{v}^{\lambda}(0)=\Phi(w_{0})\in\mathcal{M}$. While the dynamics of $\tilde{v}^{\lambda}$ might be hard to control as $\lambda\to 0$, it is easier to control the one of $\Phi(\tilde{v}^{\lambda})$. Using the chain rule, we indeed have
@@
Then using the fact that for any $w$ in a neighbourhood of $\mathcal{M}$, $D\Phi(w)\cdot\nabla F(w)=0$ (Li et al., 2021, Lemma C.2), this directly rewrites as
@@
Now note that this resembles the differential equation satisfied by $\tilde{w}^{\circ}$. The two differences being that (i) the initialisation points differ, but $\lim_{\lambda\to 0}\tilde{v}^{\lambda}(0)=\tilde{w}^{\circ}(0)$; (ii) the time derivative is on $\Phi(\tilde{v}^{\lambda})$ rather than $\tilde{v}^{\lambda}$ directly. To handle the second point, $\tilde{v}^{\lambda}$ and $\Phi(\tilde{v}^{\lambda})$ obviously converge to the same initialisation point as $\lambda\to 0$. One can then use stability of the manifold $\mathcal{M}$, thanks to Assumption 2, to show that $\sup_{t\in[0,T]}\|\tilde{v}^{\lambda}(t)-\Phi(\tilde{v}^{\lambda}(t))\|$ converges to 00 as $\lambda\to 0$. This then allows to conclude. ∎
@@
The notion of KKT points indeed extend to smooth manifolds (Bergmann and Herzog, 2019), so that under Assumption 2, the KKT points of Equation 7 are given by the points $w^{\star}\in\mathcal{M}$ satisfying $\mathrm{grad}_{\mathcal{M}}\ell_{2}(w^{\star})=0$, where we recall $\mathrm{grad}_{\mathcal{M}}\ell_{2}$ is the Riemannian gradient.
We are then able to show that $w^{\lambda}$ converges towards the set of KKT points. Note that this does not follow from Proposition 2 alone, as we also need to show that trajectory of $w^{\lambda}$ remains bounded independently of $\lambda$: we can prove this is true in our case.
@@
If Assumptions 1 and 2 hold, then for any sequence $(\lambda_{k})_{k\in\mathbb{N}}$ such that $\lambda_{k}\underset{k\to\infty}{\rightarrow}0$, the limit points of $(\lim_{t\to\infty}w^{\lambda_{k}}(t))_{k\in\mathbb{N}}$ are included in the KKT points of Equation 7.
While Proposition 3 guarantees that $w^{\lambda}$ gets arbitrarily close to KKT points of Equation 7 as $\lambda$ goes to 00, it does not imply that it has the same limit as $\tilde{w}^{\circ}$. It is however guaranteed with the additional assumption that $\tilde{w}^{\circ}(t)$ converges to a strict local minimum of the Euclidean norm on the manifold $\mathcal{M}$.
@@
Let Assumptions 1 and 2 hold and, assume additionally that $w^{\circ}(t)$ converges towards a strict local minimum $w^{\star}$ of the constrained problem (7). Then $\lim_{\lambda\to 0}\lim_{t\to\infty}w^{\lambda}(t)=w^{\star}$.
@@
When the slow limit dynamics on $\mathcal{M}$ converges towards a strict local minimum, Proposition 4 guarantees that, for small enough $\lambda$, $w^{\lambda}$ gets trapped in the vicinity of this local minimum as $t\to\infty$, allowing us to get a perfect characterisation of $\lim_{\lambda\to 0}\lim_{t\to\infty}w^{\lambda}(t)$. In particular, this double limit corresponds to the limit of the slow dynamics $\tilde{w}^{\circ}$, while the permuted limit ($\lim_{t\to\infty}\lim_{\lambda\to 0}w^{\lambda}(t)$) corresponds to the limit of the fast dynamics $w^{\mathrm{GF}}$, thanks to Proposition 1.
Comparison with Lyu et al. (2023).
@@
The work most closely related to ours is that of Lyu et al. (2023), which provides a theoretical characterization of the grokking phenomenon as a transition from the NTK regime—i.e., the unregularised flow initialised at large scales—to the rich regime, which typically converges to KKT points of Equation 7. However, their analysis does not offer a general optimisation-based perspective on the phenomenon and, in particular, does not account for the slow drift phase along the solution manifold, which we identify and characterise. Moreover, their setting is more restrictive: it assumes specific network architectures with homogeneous parameterisation and requires large initialisation scales. In contrast, our results hold outside the NTK regime and apply across a broader class of settings. In addition, their theoretical guarantees rely on taking the initialisation scale to infinity while simultaneously letting the regularisation strength tend to zero, with both rates polynomially coupled. Their analysis establishes that, for some sufficiently large time $\tilde{t}(\lambda)$, the regularised flow $w^{\lambda}$ approaches KKT points of Equation 7, but it does not provide guarantees about the asymptotic behavior beyond this time. By contrast, Proposition 3 characterises the limit points of the flow $w^{\lambda}$, offering a stronger and more complete understanding of its long-term dynamics.

### v2

@@ 4 Grokking as two-timescale dynamics
This section states our main results, where we characterise the two-timescale dynamics of the regularised gradient flow (1) in the limit $\lambda\to 0$. In Section˜4.1 we describe the first phase, the fast dynamics, where $w^{\lambda}$ approximates the unregularised gradient flow solution on finite time horizons. Section˜4.2 then identifies the second, slow dynamics happening at arbitrarily large time horizons, where $w^{\lambda}$ follows the Riemannian flow of the $\ell_{2}$ norm on the manifold $\mathcal{M}$ of stationary points.
@@ 4.1 Fast dynamics
A first simple observation is that as $\lambda\to 0$, $F_{\lambda}\to F$ uniformly on any compact of $\mathbb{R}^{d}$. From there, it seems natural that $w^{\lambda}$ should converge, at least pointwise, to $w^{\mathrm{GF}}$ as $\lambda\to 0$. A Grönwall argument indeed allows to characterise the first, fast timescale dynamics given by Proposition˜1 below.
@@
If Assumption˜1 holds, then for all $T\geq 0$, $w^{\lambda}\underset{\lambda\to 0}{\longrightarrow}w^{GF}$ in $(\mathcal{C}^{0}([0,T],\mathbb{R}^{d}),\|\cdot\|_{\infty})$
Importantly, uniform convergence only holds on finite time intervals of the form $[0,T]$, but is not true on $\mathbb{R}_{+}$. More precisely, grokking is observed when the two limits cannot be exchanged: $\lim_{\lambda\to 0^{+}}\lim_{t\to\infty}w^{\lambda}(t)\neq\lim_{t\to\infty}\lim_{\lambda\to 0^{+}}w^{\lambda}(t)=\lim_{t\to\infty}w^{\mathrm{GF}}(t)$. In Figure˜1, the endpoint of the red arrow corresponds to this first limit, while the endpoint of the blue arrow corresponds to the second. This distinction highlights a key aspect of grokking: the dynamics evolve on two different timescales. Initially, as described by Proposition˜1, the regularised flow tracks the unregularised one. But at much larger time horizons, the regularised dynamics begin to diverge.
@@ 4.2 Slow dynamics
$\phi(w,t)=w-\int_{0}^{t}\nabla F(\phi(w,s))\ \mathrm{d}s,\qquad\forall(w,t)\in\mathbb{R}^{d}\times\mathbb{R}_{+}.$
Note that $\phi(w,t)$ simply corresponds to the solution of the gradient flow of Equation˜2 at time $t$ when initialised in $w$. When possible to define—i.e., when the gradient flow admits a limit point in $\mathbb{R}^{d}$—we define the mapping $\Phi$ as
@@
Thanks to Assumption˜1, the unregularised flow initialised in $w_{0}$ admits the limit point $\Phi(w_{0})$, which is necessarily a stationary point of the loss, i.e., $\nabla F(\Phi(w_{0}))=0$. Assumption˜2 then ensure that the mapping $\Phi$ is defined and $\mathcal{C}^{2}$ on some neighbourhood of $\mathcal{M}$, thanks to a result of falconer1983differentiation.
@@
If Assumptions˜1 and 2 hold, there exists an open neighbourhood $U$ of $\mathcal{M}$ such that $\Phi$ is defined and $\mathcal{C}^{2}$ on $U$.
@@
Our goal in this section is to study the limit function $\lim_{\lambda\to 0}\tilde{w}^{\lambda}$. Intuitively, the $\frac{1}{\lambda}\nabla F$ term in Equation˜5 will enforce this limit to stay on the stationary manifold $\mathcal{M}$ for any $t>0$. The slow dynamics will be shown to approximate the Riemannian flow of the squared Euclidean norm on the stationary manifold $\mathcal{M}$ for $t>0$. This limit flow is defined by $\tilde{w}^{\circ}$, which is the solution of the following differential equation on $\mathbb{R}_{+}$, for the function $\ell_{2}:w\mapsto\|w\|_{2}^{2}/2$,
@@
Recall that the Riemannian gradient is $\mathrm{grad}_{\mathcal{M}}\ \ell_{2}(w)=P_{\mathrm{Ker}(\nabla^{2}F(w))}(w)$ for any $w\in\mathcal{M}$. Denoting $D\Phi_{w}$ the differential of $\Phi$ at $w$, li2021happens proved that for any $w\in\mathcal{M}$, $P_{\mathrm{Ker}(\nabla^{2}F(w))}=D\Phi_{w}$, i.e., the differential of $\Phi$ at $w$ is given by the orthogonal projection onto the kernel space of the Hessian of $F$. In consequence, $\tilde{w}^{\circ}$ also satisfies the following differential equation:
@@
Using this alternative description of $\tilde{w}^{\circ}$, we can now prove our main result, given by Proposition˜2.
@@
If Assumptions˜1 and 2 hold, then for all $T,\varepsilon>0$, we have $\tilde{w}^{\lambda}\underset{\lambda\to 0}{\longrightarrow}\tilde{w}^{\circ}$ in $(\mathcal{C}^{0}([\varepsilon,T],\mathbb{R}^{d}),\|\cdot\|_{\infty})$, where $\tilde{w}^{\circ}$ is the unique solution on $\mathbb{R}_{+}$ of the differential equation (6).
Proposition˜2 states that the slow dynamics $\tilde{w}^{\lambda}$ converges uniformly to $\tilde{w}^{\circ}$ as $\lambda\to 0$ on any compact interval of the form $\varepsilon,T]$. Note that excluding 0 from this interval (i.e., $\varepsilon>0$) is necessary. Indeed, uniform convergence cannot happen on an interval of the form $(0,T]$, since $\tilde{w}^{\lambda}(0)=w_{0}$ for any $\lambda>0$ and $\tilde{w}^{\circ}(0)=\Phi(w_{0})$. In particular, Proposition˜2 leads to pointwise convergence of $\tilde{w}^{\lambda}$: we have
@@
This limit function $\lim_{\lambda\to 0}\tilde{w}^{\lambda}$ is non-continuous at 0. Indeed the whole fast dynamics, which follows the unregularised flow, happens at that 0 point in the limit $\lambda\to 0$. On the other hand, Proposition˜2 describes the second phase of the dynamics, starting from the convergence point of the unregularized flow $\Phi(w_{0})$ – at the rescaled time $0^{+}$ – and following the Riemannian flow on $\mathcal{M}$.
Note that, once the junction between the slow and fast dynamics is carefully handled via Lemma˜2 in Section˜A.2, Proposition˜2 can be derived from fatkullin2010reduced, which heavily relies on the technical result of katzenberger1990solutions. However, for the sake of completeness and readability, we provide a concise and self-contained proof of Proposition˜2, avoiding the use of heavyweight methods from katzenberger1990solutions and relying on weaker assumptions.
@@
We here provide a sketch of proof with the key arguments leading to Proposition˜2. Its complete and detailed proof can be found in Section˜A.3. We first define the shifted slow dynamics $\tilde{v}^{\lambda}$ for any $t\geq 0$ as $\tilde{v}^{\lambda}(t)=\tilde{w}^{\lambda}(t+t(\lambda))$, where $t(\lambda)$ is the “junction point” between the two dynamics given by Lemma˜2 in Section˜A.2, and satisfies $\lim_{\lambda\rightarrow 0}t(\lambda)=0$. Using Lemma˜2, $\tilde{v}^{\lambda}$ then follows the same differential equation as $\tilde{w}^{\lambda}$, with an initial condition now satisfying $\lim_{\lambda\to 0}\tilde{v}^{\lambda}(0)=\Phi(w_{0})\in\mathcal{M}$. While the dynamics of $\tilde{v}^{\lambda}$ might be hard to control as $\lambda\to 0$, it is easier to control the one of $\Phi(\tilde{v}^{\lambda})$. Using the chain rule, we indeed have
@@
Then using the fact that for any $w$ in a neighbourhood of $\mathcal{M}$, $D\Phi(w)\cdot\nabla F(w)=0$ (li2021happens, Lemma C.2), this directly rewrites as
@@
Now note that this resembles the differential equation satisfied by $\tilde{w}^{\circ}$. The two differences being that (i) the initialisation points differ, but $\lim_{\lambda\to 0}\tilde{v}^{\lambda}(0)=\tilde{w}^{\circ}(0)$; (ii) the time derivative is on $\Phi(\tilde{v}^{\lambda})$ rather than $\tilde{v}^{\lambda}$ directly. To handle the second point, $\tilde{v}^{\lambda}$ and $\Phi(\tilde{v}^{\lambda})$ obviously converge to the same initialisation point as $\lambda\to 0$. One can then use stability of the manifold $\mathcal{M}$, thanks to Assumption˜2, to show that $\sup_{t\in[0,T]}\|\tilde{v}^{\lambda}(t)-\Phi(\tilde{v}^{\lambda}(t))\|$ converges to 0 as $\lambda\to 0$. This then allows to conclude. ∎
@@
The notion of KKT points indeed extend to smooth manifolds (bergmann2019intrinsic), so that under Assumption˜2, the KKT points of Equation˜7 are given by the points $w^{\star}\in\mathcal{M}$ satisfying $\mathrm{grad}_{\mathcal{M}}\ell_{2}(w^{\star})=0$, where we recall $\mathrm{grad}_{\mathcal{M}}\ell_{2}$ is the Riemannian gradient.
We are then able to show that $w^{\lambda}$ converges towards the set of KKT points. Note that this does not follow from Proposition˜2 alone, as we also need to show that trajectory of $w^{\lambda}$ remains bounded independently of $\lambda$: we can prove this is true in our case.
@@
If Assumptions˜1 and 2 hold, then for any sequence $(\lambda_{k})_{k\in\mathbb{N}}$ such that $\lambda_{k}\underset{k\to\infty}{\rightarrow}0$, the limit points of $(\lim_{t\to\infty}w^{\lambda_{k}}(t))_{k\in\mathbb{N}}$ are included in the KKT points of Equation˜7.
While Proposition˜3 guarantees that $w^{\lambda}$ gets arbitrarily close to KKT points of Equation˜7 as $\lambda$ goes to 0, it does not imply that it has the same limit as $\tilde{w}^{\circ}$. It is however guaranteed with the additional assumption that $\tilde{w}^{\circ}(t)$ converges to a strict local minimum of the Euclidean norm on the manifold $\mathcal{M}$.
@@
Let Assumptions˜1 and 2 hold and, assume additionally that $w^{\circ}(t)$ converges towards a strict local minimum $w^{\star}$ of the constrained problem (7). Then $\lim_{\lambda\to 0}\lim_{t\to\infty}w^{\lambda}(t)=w^{\star}$.
@@
When the slow limit dynamics on $\mathcal{M}$ converges towards a strict local minimum, Proposition˜4 guarantees that, for small enough $\lambda$, $w^{\lambda}$ gets trapped in the vicinity of this local minimum as $t\to\infty$, allowing us to get a perfect characterisation of $\lim_{\lambda\to 0}\lim_{t\to\infty}w^{\lambda}(t)$. In particular, this double limit corresponds to the limit of the slow dynamics $\tilde{w}^{\circ}$, while the permuted limit ($\lim_{t\to\infty}\lim_{\lambda\to 0}w^{\lambda}(t)$) corresponds to the limit of the fast dynamics $w^{\mathrm{GF}}$, thanks to Proposition˜1.
Role of initialisation scale.
The initialisation scale strongly influences the behavior of the unregularised flow and, consequently, the first phase of training under weight decay. This dependence on scale is well-documented in the literature (chizat2019lazy). While a complete theoretical understanding remains open, it is widely accepted that small initialisation scales correspond to the rich regime, in which implicit bias drives the model toward interpolating solutions with smaller weight norms, typically associated with better generalization. In contrast, large initialisation scales give rise to the lazy or NTK regime, where features change little during training. This regime behaves similarly to random feature models and tends to produce interpolators with weaker generalisation performance.
In our framework, this distinction has a direct consequence. With a small initialisation scale, the point $\Phi(w^{\lambda}(0))$ reached after the first phase already exhibits a small norm–possibly corresponding to a KKT point of Equation˜7–so no second grokking phase occurs, as the system has effectively converged. Conversely, with a large initialisation scale, $\Phi(w^{\lambda}(0))$ retains a large norm, which triggers substantial movement during the second, grokking phase.
The grokking transition is not sudden!
In the literature, grokking is often described as a “sudden drop” in the validation loss following an extended phase of overfitting. We argue here that this drop only seems sudden when training time is plotted on a logarithmic scale, whereas in fact it unfolds over a characteristic duration of order $1/\lambda$, following a plateau of comparable duration $1/\lambda$. Indeed, Proposition˜2 predicts that the drop takes place within the time interval $\varepsilon/\lambda,M/\lambda]$, where $\varepsilon$ is a small time independent of $\lambda$ (think of $\varepsilon$ as the time it takes for $\tilde{w}^{\circ}$ to move very slightly away from $w^{\rm GF}_{\infty}$), and $M$ corresponds to the typical time required for $\tilde{w}^{\circ}(t)$ to approach its limit $\lim_{t\to\infty}\tilde{w}^{\circ}(t)$. Prior to this interval, the parameters evolve according to the unregularised flow $w^{\rm GF}(t)$. Let $M^{\prime}$ denote the typical time it takes for $w^{\rm GF}(t)$ to reach $w^{\rm GF}_{\infty}$; then the plateau extends over the interval $M^{\prime},\varepsilon/\lambda]$. As a result, on a logarithmic time scale, the drop occurs within an interval of length $\ln(M/\varepsilon)$, while the preceding plateau spans roughly $\ln(\varepsilon/\lambda)$ on the same scale. This explains why, as $\lambda\to 0$, the drop appears abrupt compared to the plateau in log-scale plots. In contrast, when viewed in linear time, the drop actually extends over a duration comparable to that of the preceding plateau, producing a markedly different visual impression.
Comparison with lyu2023dichotomy.
@@
The work most closely related to ours is that of lyu2023dichotomy, which provides a theoretical characterization of the grokking phenomenon as a transition from the NTK regime—i.e., the unregularised flow initialised at large scales—to the rich regime, which typically converges to KKT points of Equation˜7. However, their analysis does not offer a general optimisation-based perspective on the phenomenon and, in particular, does not account for the slow drift phase along the solution manifold, which we identify and characterise. Moreover, their setting is more restrictive: it assumes specific network architectures with homogeneous parameterisation and requires large initialisation scales. In contrast, our results hold outside the NTK regime and apply across a broader class of settings. In addition, their theoretical guarantees rely on taking the initialisation scale to infinity while simultaneously letting the regularisation strength tend to zero, with both rates polynomially coupled. Their analysis establishes that, for some sufficiently large time $\tilde{t}(\lambda)$, the regularised flow $w^{\lambda}$ approaches KKT points of Equation˜7, but it does not provide guarantees about the asymptotic behavior beyond this time. By contrast, Proposition˜3 characterises the limit points of the flow $w^{\lambda}$, offering a stronger and more complete understanding of its long-term dynamics.